Circular No. 7122

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1999am IN CGCG 60-009 G. Aldering and P. Nugent, Lawrence Berkeley National Laboratory,

on behalf of the Supernova Cosmology Project (SCP), report the discovery of a supernova (unfiltered mag ∼ 17.8) on Feb. 18 UT as part of a joint collaboration between the SCP and E. Helin, S. Pravdo, D. Rabinowitz, and K. Lawrence (Near-Earth Asteroid Tracking project, Jet Propulsion Laboratory). SN 1999am is located at α = 8^{h}24^{m}10^{s}.88, δ = +12^{o}06′58′′.6 (equinox 2000.0), which is 11′′ west of the center of CGCG 060-009. W. Kunkel (Carnegie Observatories) and M. Phillips (Carnegie Institution of Washington) report that a spectrum of the new object taken on Mar. 3 with the 2.54-m reflector at Las Campanas shows it to be a type-Ia supernova, nearly 3 weeks after maximum light at a redshift of ∼ 12 000 km/s.

SUPERNOVA 1999ac IN NGC 6063 M. M. Phillips, Las Campanas Observatory, Carnegie Institution of

Washington, writes: “A CCD spectrum (range 365–930 nm) obtained by M. M. Phillips and W. Kunkel, Carnegie Observatories, on Feb. 28 UT with the Las Campanas 2.5-m duPont telescope shows that SN 1999ac is a peculiar, 1991T-like type-Ia event that has been caught near, or a few days before, maximum. This classification is confirmed by A. V. Filippenko, University of California at Berkeley, who notes that the spectrum is similar, but somewhat latter in epoch, to the spectrum of SN 1999aa described on IAUC 7108. Like SN 1991T, the Fe III lines at 430 and 500 nm in the spectrum of SN 1999ac are observed to be strong and the Si II 635.5-nm absorption weak; but unlike 1991T, the Ca II H and K aborption in SN 1999ac is strong and well-developed.”

Unfiltered CCD magnitude estimates: Feb. 28.111 UT, 15.3 ± 0.3 (L. Kiss and K. Sárneczky, Szeged, Hungary); Mar. 3.726, 15.7 (S. Yoshida and K. Kadota, Ageo City, Japan).

SUPERNOVAE 1999X AND 1999aa Unfiltered CCD estimates by S. Yoshida and K. Kadota, Ageo City,

Japan: SN 1999X in CGCG 180-22, Feb. 7.641 UT, 16.5; 22.710, 16.8. SN 1999aa in NGC 2595, Mar. 3.693, 15.8.

1999 March 6  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green